OpenStudy (anonymous):

I really dont get this one help? x^2-x-6 / 2x^2+x-6

OpenStudy (anonymous):

factor the top and the bottom. do you know how to factor?

OpenStudy (anonymous):

somewhat

OpenStudy (anonymous):

alright what would the numerator be factored?

OpenStudy (anonymous):

you need two numbers that add to -1 and multipy to get -6

OpenStudy (anonymous):

\[x^2-x-6\] \[?*?=-6,?+?=-1\] there is another way of doing this if you're stumped. it's more systematic and takes longer, but helps you before you can do coefficient 1 quadratic equations.

OpenStudy (anonymous):

would you like me to show you the other way?

OpenStudy (anonymous):

yeah please

OpenStudy (anonymous):

alright so first take the coefficient of the term \[x^2\] the coefficient is 1. next take the constant (the number without x. That number is -6.multiply these numbers. \[1*-6=-6\] figure out all the possible combinations to multiply -6 \[1*-6,2*-3,3*-2,6*-1\] now take these and see which add up to -1 \[1+-6=-5\] \[2+(-3)=-1\] as you can see 2 and -3 are the numbers we're looking for. Now take the 2 numbers, multiply them by x and replace the -x in your equation by them. you should get this \[x^2+2x-3x-6\] now group the first two and the second two \[(x^2+2x)+(-3x-6)\] in each group you can see a common factor for the first it is x for the second it is -3 doing this gets you \[x(x+2)-3(x+2)\] when you se x+2 in both parenthesis you know you have done correctly. You only need one x+2 and the outside part \[(x+2)(x-3)\] this is the same as your original \[x^2-1x-6\]

OpenStudy (anonymous):

so now that you know the top is (x+2)(x-3), think youcan do the bottom? and tell me what you ge

OpenStudy (anonymous):

okay let me try it

OpenStudy (anonymous):

is it something like 2x-3 ?

OpenStudy (anonymous):

yes and the other part of the factor is?

OpenStudy (anonymous):

(2x-3)(x+?)

OpenStudy (anonymous):

the top part?

OpenStudy (anonymous):

when you factored the bottom you grouped, you got something like this correct? 2x(x+?)-3(x+?)

OpenStudy (anonymous):

2x-3 is part of the factor but you still needthe (x+?)

OpenStudy (anonymous):

you should have got (x+2)(2x-3) when you put the numerator and denominator together you get \[\frac{(x+2)(x-3)}{(2x-3)(x+2)}=\frac{x-3}{2x-3}\]